OpenStudy (sasogeek):

a particle of mass 2kg is acted on by two forces of magnitude 4N and 2N in the direction i + 2j and 2i - j respectively. Calculate the magnitude of the acceleration of the particle

OpenStudy (anonymous):

Net force in x direction = 8N

OpenStudy (anonymous):

Net force in y direction = 6N

OpenStudy (anonymous):

Magnitude=10 by a^2=b^2+c^2

OpenStudy (jamesj):

So the unit vector in direction i + 2j is v = 1/sqrt(5)(i + 2j) Find the unit vector for direction (2i-j), call it w, say Then the total force is 4v + 2w

OpenStudy (anonymous):

a=F/m =10/2 = 5m/s^2

OpenStudy (jamesj):

@EE: what you've written for the force above is incorrect, because you didn't normalize the direction vectors.

OpenStudy (anonymous):

F = 4(i+2j) + 2(2i-j) = (8i+6j)

OpenStudy (jamesj):

No. Because i + 2j gives the direction. But that vector has magnitude 5.

OpenStudy (anonymous):

Yeah it's probably wrong.

OpenStudy (jamesj):

so you can't just blithely multiply

OpenStudy (sasogeek):

thanks a lot :)

OpenStudy (jamesj):

(make that: the vector i + 2j has magnitude sqrt(5) )